OpenStudy (anonymous):

-3/11 divided by 6/11

OpenStudy (immanuelv):

do you have a calculator?

OpenStudy (anonymous):

no i don't

OpenStudy (immanuelv):

well since they both have the same base you can divide (-3/6) and thats your answer

OpenStudy (anonymous):

thank you

OpenStudy (immanuelv):

you're welcome

OpenStudy (anonymous):

so it's 2 but is it regular or with the minus on it

OpenStudy (anonymous):

-3/11/6/11=-3/6=-1/2

OpenStudy (immanuelv):

-1/2

TheSmartOne (thesmartone):

\[-\frac{ 3 }{ 11 }\div \frac{ 6 }{ 11 }=-\frac{ 3 }{ 11 }\times \frac{ 11 }{6 }=\frac{ -3 }{ 6 }=\frac{ -1 }{ 2 }\]

TheSmartOne (thesmartone):

Does this explain it @momszoo

OpenStudy (anonymous):

no i don't have a calculator says the person who is on a computer?

TheSmartOne (thesmartone):

so true ^^^